# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

### v2

@@ 6.5. Proof of Theorem F
and it follows that there are two $\operatorname{Pin}^{+}$-structures $(A(2p),\pm\phi_{A(2p)})$ and two $\operatorname{Pin}^{+}$-structures $(B(2p),\pm\phi_{B(2p)})$ [35, Corollary 6.4], . Moreover, the $\operatorname{Pin}^{+}$-structures on the same underlying smooth four-manifold are mutual inverses in the fourth $\operatorname{Pin}^{+}$-cobordism group $\Omega^{\operatorname{Pin}^{+}}_{4}$ and the values
@@
belong to different classes in $\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16$ for any $k\in\mathbb{Z}_{>0}$ and that the four-manifolds $A(2p)$ and $B(2p)$ are not stably-diffeomorphic. This settles the claim of Item (5). Item (6) follows from the third clause of Theorem 28 and the corresponding coverings of the decompositions (6.7) and (6.8).
Finally, to construct the closed non-smoothable topological four-manifold $\ast A(2p)$, we substitute $\mathbb{R}P^{4}$ with the closed non-smootable four-manifold $\ast\mathbb{R}P^{4}$ that is homotopy equivalent to the real projective four-space in (6.7). The latter four-manifold exists by a computation using the surgery exact sequence [39, Chapter 14] and an explicit construction of it was given by Ruberman in [32]. The existence and uniqueness results of a topological tubular neighborhood $\nu(\alpha_{\ast\mathbb{R}P^{4}})$ for a locally flat simple loop $\alpha_{\ast\mathbb{R}P^{4}}\subset\ast\mathbb{R}P^{4}$ are found in [16, Section 5.1]. Choose a loop that generates the fundamental group and assemble
@@
The existence of a homotopy equivalence $\ast\mathbb{R}P^{4}\rightarrow\mathbb{R}P^{4}$ implies that there is a homotopy equivalence $\ast A(2p)\rightarrow A(2p)$. From (6.11) and the related cobordism, we have that the Kirby-Siebenmann invariant is $\operatorname{KS}(\ast A(2p))=\operatorname{KS}(\ast\mathbb{R}P^{4})=1\in\mathbb{Z}/2$ [16, Theorem 9.2] (see [16, Section 9.2] for a description of the construction of this invariant). Thus, the four-manifold $\ast A(2p)$ is non-smoothable for any odd $p\in\mathbb{Z}_{>0}$. $\square$

### v3

@@ 6.5. Proof of Theorem F
and it follows that there are two $\operatorname{Pin}^{+}$-structures $(A(2p),\pm\phi_{A(2p)})$ and two $\operatorname{Pin}^{+}$-structures $(B(2p),\pm\phi_{B(2p)})$ [41, Corollary 6.4], . Moreover, the $\operatorname{Pin}^{+}$-structures on the same underlying four-manifold are mutual inverses in the fourth $\operatorname{Pin}^{+}$-cobordism group $\Omega^{\operatorname{Pin}^{+}}_{4}$ and the values
@@
belong to different classes in $\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16$ for any $k\in\mathbb{Z}_{>0}$ and that the four-manifolds $A(2p)$ and $B(2p)$ are not stably-diffeomorphic. This settles the claim of Item (5). Item (6) follows from the third clause of Theorem 32 and the corresponding covers of the decompositions (6.18) and (6.19).
Finally, to construct the closed non-smoothable topological four-manifold $\ast A(2p)$, we substitute ${\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$ with the closed non-smootable four-manifold $\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$ that is homotopy equivalent to the real projective four-space in (6.18). The latter four-manifold exists by a computation using the surgery exact sequence [47, Chapter 14] and an explicit construction of it was given by Ruberman in [38]. The existence and uniqueness results of a topological tubular neighborhood $\nu(\alpha_{\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}})$ for a locally flat simple loop $\alpha_{\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}}\subset\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$ are found in [17, Section 5.1]. Choose a loop that generates the fundamental group and assemble
@@
The existence of a homotopy equivalence $\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}\rightarrow{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$ implies that there is a homotopy equivalence $\ast A(2p)\rightarrow A(2p)$. From (6.22) and the related cobordism, we have that the Kirby-Siebenmann invariant is $\operatorname{KS}(\ast A(2p))=\operatorname{KS}(\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4})=1\in\mathbb{Z}/2$ [17, Theorem 9.2] (see [17, Section 9.2] for a description of the construction of this invariant). Thus, the four-manifold $\ast A(2p)$ is non-smoothable for any odd $p\in\mathbb{Z}_{>0}$. The homeomorphism class of its universal cover is pinned down by applying Freedman’s theorem [21, Theorem 1.2.27]. $\square$